OpenStudy (anonymous):

Solve these quadradics by factoring; x^2 = 14x – 49. x^2 + 18 = –9x. 4x^2 = 12x + 40.

OpenStudy (anonymous):

1. (x-7)^2

jhonyy9 (jhonyy9):

1. (x-7)(x-7) 2. (x+6)(x+3) 3.4(x2 -3x-10)=4(x+2)(x-5) so for 1. x=7 for 2. x=-6 and x=-3 for 3. x=-2 and x=5